OpenStudy (anonymous):

What's 2 1/5+1 1/10? A. 3 1/10 B. 3 1/5 C. 3 3/10 D. 3 3/5

OpenStudy (anonymous):

I think it's C

OpenStudy (anonymous):

You're correct... do you want to go over it or are you okay?

OpenStudy (anonymous):

I want to go over it please

OpenStudy (anonymous):

Okay, so before you can add the fractions 1/5 and 1/10, they need to have a common denominator... The lowest common multiple of 5 and 10 is 10, so you need to change 1/5 to x/10... You do that by multiplying the numerator and the denominator by 2. 1X2=2 and 5X2=10... So now you have 2/10 + 1/10... add the numerators, and you have 3/10, and I'm sure you know to just add the whole numbers that are in the equation..2+1=3... so 3 and 3/10

OpenStudy (timaashorty):

For now ignore the 2 and one. Focus on this: \[\frac{ 1 }{ 5 } + \frac{ 1 }{ 10 }\] Change both the denominators to make them look the same. By doing that you need to find the least common multiple for both. I will convert them to both denometers to 10's. \[1/5 x2 = \frac{ 2 }{ 10 } so, \frac{ 2 }{ 10 } + \frac{ 1 }{ 10 } = 3/10\] \[add the whole numbers, 2 + 1 = 3\] So therefore, it is \[3 \frac{ 3 }{ 10}\]

OpenStudy (anonymous):

I understand it a little more! Thanks

OpenStudy (timaashorty):

You're welcome, If need further help just let us know [: